OpenStudy (anonymous):

Carly is laying carpet in a square foyer with sides of 5.67 ft each. Estimate the amount of carpet she will need for the foyer, using whole numbers.

OpenStudy (tyteen4a03):

The area of a square can be found using this formula: \(s^2\), where s is one side of the square. Also, \(s^2 = s*s, s^3 = s*s*s, s^4 = s*s*s*s\), and so on.